# Refining Neural Activation Patterns for Layer-Level Concept Discovery in Neural Network-Based Receivers

### v1

@@ G.3 NUCPI Analysis of User Equipment (UE) Speed
Appendix H Cluster Statistics and DBCV Scores for the Example Case
This appendix presents detailed cluster statistics and DBCV scores for the example case shown in Figure 7. The baseline method results are shown in Listings 1 and 3, while the proposed method results are in Listings 2 and 4.
Cluster #0, 7 samples
Persistence score: 0.01
Data probability: 1.00 +- 0.01
SNR: 7.60 +- 11.97 dB
UE speed: 71.76 +- 31.59 km/h
Delay spread: 111.52 +- 72.39 ns
Channel model: 0.43 +- 0.49
CDL-A (nlos): 7 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 0 samples
BER: 0.0669 +- 0.0746
SER: 0.3297 +- 0.3387
Cluster #1, 38373 samples
Persistence score: 0.12
Data probability: 0.98 +- 0.05
SNR: 7.50 +- 10.10 dB
UE speed: 62.08 +- 35.99 km/h
Delay spread: 155.57 +- 83.73 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 18686 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 19687 samples
BER: 0.0569 +- 0.0649
SER: 0.3004 +- 0.3065
Cluster #2, 6 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 10.09 +- 9.24 dB
UE speed: 57.99 +- 24.47 km/h
Delay spread: 155.65 +- 89.27 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 6 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 0 samples
BER: 0.0402 +- 0.0383
SER: 0.2617 +- 0.2329
Listing 1: Cluster statistics for baseline methodology in example case in Figure 7
Cluster #0, 14 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 6.52 +- 11.25 dB
UE speed: 47.24 +- 32.21 km/h
Delay spread: 144.79 +- 75.15 ns
Channel model: 0.71 +- 0.45
CDL-A (nlos): 8 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 6 samples
BER: 0.0643 +- 0.0719
SER: 0.3190 +- 0.3241
Cluster #1, 5 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 10.01 +- 9.96 dB
UE speed: 90.18 +- 25.18 km/h
Delay spread: 221.67 +- 58.79 ns
Channel model: 0.40 +- 0.49
CDL-A (nlos): 2 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 3 samples
BER: 0.0496 +- 0.0609
SER: 0.2647 +- 0.3241
Cluster #2, 5 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 4.45 +- 9.56 dB
UE speed: 63.30 +- 16.96 km/h
Delay spread: 198.46 +- 30.81 ns
Channel model: 0.60 +- 0.49
CDL-A (nlos): 3 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 2 samples
BER: 0.0745 +- 0.0674
SER: 0.3771 +- 0.3149
Cluster #3, 6 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 10.41 +- 10.36 dB
UE speed: 68.90 +- 43.58 km/h
Delay spread: 119.16 +- 79.94 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 2 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 4 samples
BER: 0.0485 +- 0.0584
SER: 0.2651 +- 0.3016
Cluster #4, 10 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 11.19 +- 8.86 dB
UE speed: 76.13 +- 37.53 km/h
Delay spread: 157.09 +- 72.29 ns
Channel model: 0.80 +- 0.40
CDL-A (nlos): 4 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 6 samples
BER: 0.0346 +- 0.0577
SER: 0.1866 +- 0.2873
Cluster #5, 6 samples
Persistence score: 0.01
Data probability: 1.00 +- 0.00
SNR: 8.93 +- 7.36 dB
UE speed: 42.42 +- 29.41 km/h
Delay spread: 185.68 +- 97.80 ns
Channel model: 0.67 +- 0.47
CDL-A (nlos): 4 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 2 samples
BER: 0.0360 +- 0.0425
SER: 0.2275 +- 0.2441
Cluster #6, 33063 samples
Persistence score: 0.07
Data probability: 1.00 +- 0.01
SNR: 7.45 +- 10.11 dB
UE speed: 61.97 +- 36.01 km/h
Delay spread: 155.05 +- 83.93 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 16415 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 16648 samples
BER: 0.0572 +- 0.0651
SER: 0.3021 +- 0.3070
Cluster #7, 6 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 4.10 +- 10.52 dB
UE speed: 55.36 +- 27.38 km/h
Delay spread: 139.33 +- 69.94 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 2 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 4 samples
BER: 0.0837 +- 0.0712
SER: 0.4219 +- 0.3209
Listing 2: Cluster statistics for proposed methodology in example case in Figure 7
DBCV for clustering: 0.0555
DBCV for cluster #0: 0.1843
DBCV for cluster #1: 0.0578
DBCV for cluster #2: 0.2587
Listing 3: DBCV scores for baseline methodology in example case in Figure 7
DBCV for clustering: 0.0141
DBCV for cluster #0: 0.0244
DBCV for cluster #1: 0.0311
DBCV for cluster #2: 0.0752
DBCV for cluster #3: 0.0685
DBCV for cluster #4: 0.1635
DBCV for cluster #5: 0.1868
DBCV for cluster #6: 0.0170
DBCV for cluster #7: 0.0279
Listing 4: DBCV scores for proposed methodology in example case in Figure 7
Appendix I On the Illusory Advantages of UMAP Dimensionality Reduction on the OUT-POST Layer of the DeepRx Model
While UMAP is commonly employed to enhance the separability of high-dimensional data in lower-dimensional spaces, our experiments indicate that its application to OUT-POST layer activations in the DeepRx model may introduce distortions that compromise interpretability. In particular, we suspect that UMAP may “tear” the embedding space, leading to artificially induced clustering that lacks correspondence to meaningful channel parameter variations. Figures 32 and 32 illustrate these divergent clustering outcomes. Although UMAP superficially increases cluster compactness, it appears to do so at the cost of semantic coherence—highlighting the need for caution when applying nonlinear dimensionality reduction in domains where physical interpretability is critical.
This suspected tearing behavior aligns with known limitations of UMAP. The algorithm prioritizes local neighborhood preservation under assumptions of uniform data distribution and local manifold connectivity. These assumptions can result in global structural distortions and discontinuities in the embedded space (McInnes et al. (2018)). Coenen et al. (2019) provide an intuitive visualization of this effect, showing how UMAP may artificially separate continuous regions. Damrich and Hamprecht (2021) further offer a theoretical analysis, attributing these artifacts to the optimization objective and the underlying graph construction. These insights are consistent with our observations and underscore the importance of hyperparameter tuning and critical interpretation of UMAP results—particularly in contexts where maintaining topological and semantic continuity is essential.
Figure 31: Clustering result after UMAP reduction to five dimensions. Although over half the data (24k samples) is assigned to clusters, the resulting 1,176 clusters show no clear physical interpretation. Clustering was based on EoM, minimum cluster size of 5, and proposed methodology.
Figure 32: Clustering result without UMAP. A bit fewer samples (19k samples) are clustered, but the 139 resulting clusters are more interpretable (clusters on low and high SNR ends). Clustering was based on EoM, minimum cluster size of 5, and proposed methodology.
Appendix J Examples of 2D UMAP Embeddings Using Unaggregated and Unnormalized DeepRx Activations
In this appendix, we present two-dimensional UMAP embeddings of the raw (i.e., unaggregated and unnormalized) activations from the DeepRx model (Appendix D) for the B3-POST and OUT-POST layers. As discussed in the main paper, direct clustering of these activations using HDBSCAN was unsuccessful, with all samples classified as outliers—likely due to the high dimensionality and lack of preprocessing in the feature space.
Additionally, an HDBSCAN run on the B1-PRE layer, using a single parameter configuration, failed to complete after approximately two weeks (341 hours) of execution. The exact cause of termination remains unclear, but the duration and lack of output suggest a computational bottleneck or instability in the clustering process.
While the use of UMAP occasionally enabled some clustering structure to emerge, the results were highly sensitive to both the layer and the chosen parameters, and no consistent patterns were observed. However, despite these limitations, the 2D UMAP projections of these raw activations reveal a continuous low-dimensional manifold primarily structured by SNR. This trend is visible in both B3-POST and OUT-POST layers (see Figure 34 and Figure 38) and aligns with the structure previously observed in preprocessed activations using both the baseline and proposed methodologies (Figure 7).
Figure 34 suggests that, for layer B3-POST, lower UE speeds tend to concentrate near the center of the manifold, while higher speeds deviate outward. However, this observation is based solely on visual inspection and exhibits considerable noise—some higher speed samples also appear near the center. A more rigorous, quantitative analysis would be necessary to draw definitive conclusions. This trend is not observed in the OUT-POST layer (Figure 38).
Figure 40 shows that some CDL-A samples in the OUT-POST layer deviate slightly from the main manifold, though this effect is subtle and not evident in the corresponding B3-POST visualization (Figure 36).
Finally, we observe no clear visual structure related to delay spread in either B3-POST (Figure 36) or OUT-POST (Figure 40) layers.
Figure 33: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by Signal-to-Noise Ratio (SNR). The manifold structure suggests an organization influenced by SNR.
Figure 34: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by User Equipment (UE) speed. A tendency for lower speeds to concentrate near the center is observable, though with noise.
Figure 35: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by delay spread values. No clear structure related to delay spread is visually apparent.
Figure 36: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by channel model. No distinct separation between different channel models is observed.
Figure 37: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by Signal-to-Noise Ratio (SNR). Similar to B3-POST, a clear SNR-driven structure is evident.
Figure 38: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by User Equipment (UE) speed. No consistent pattern associated with UE speed is observed.
Figure 39: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by delay spread values. Delay spread does not appear to influence the structure of the embedding.
Figure 40: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by channel model class. Some CDL-A samples show minor deviation from the main manifold.
Appendix K Compute Resources and Environment
We outline the compute resources, memory and storage specifications, software stack, and execution time for our experimental setup. The execution time figures are approximate, derived from logs, rather than precise measurements from instrumentation or formal profiling tools.
K.1 Compute Environment
All experiments were conducted using the following cloud-based compute resources:
  • GPU: Multi-Instance GPU (MIG) instances, each providing 10GB of dedicated GPU memory. The performance of each instance is approximately equivalent to a lower-end dedicated GPU such as the NVIDIA T4, or roughly one-seventh (1/7th) of a high-end data center GPU (NVIDIA A100 80GB).
  • CPU: 16 virtual CPU cores. These cores are comparable to those from 3rd Generation Intel Xeon Scalable processors (Ice Lake architecture), commonly found in enterprise and data center environments.
  • Memory: 512 GB system RAM
  • Storage: 1 TB HDD
  • Software Stack: Python 3.10, PyTorch 2.6, TensorFlow 2.6, and many other dependencies. The most essential dependencies are specified in Listing 5.
h5py==3.13.0
hdbscan==0.8.40
joblib==1.4.2
matplotlib==3.10.1
numba==0.61.0
numpy==2.1.3
pandas==2.2.3
scikit-learn==1.6.1
scipy==1.15.2
sionna==0.10.0
torch==2.6.0
torchvision==0.21.0
umap-learn==0.5.7
tensorflow==2.6.4
Listing 5: Most essential Python requirements used in requirements.txt format.
[…]

### v2

@@ C.3.1 NUCPI Analysis of Delay Spread
C.3.2 NUCPI Analysis of User Equipment (UE) Speed
We also applied NUCPI analysis to UE speed, evaluating layerwise explainer performance for two layers. As with delay spread (see Appendix C.3.1), the explainer models exhibit consistent overfitting: training losses decrease, but evaluation losses do not improve, as shown in Figures 31 and 32. Based on this behavior, we conclude that DeepRx’s internal activations likely do not encode UE speed in a directly predictive way.
In traditional systems, UE speed—or Doppler spread more technically—is estimated by aggregating over tens of slots (yucek2005; zhao2009). In contrast, DeepRx operates on a single slot without a memory of past slots. Despite this, DeepRx can still handle frequency offset effects caused by user movement (and other factors). A similar explanation as with delay spread may apply: the model may not need to explicitly estimate frequency offset or Doppler spread to effectively compensate for them.
Appendix D Uniform Manifold Approximation and Projection (UMAP) Methodology for Manifold Analysis
We use Uniform Manifold Approximation and Projection (UMAP) mcinnes2018umap to visualize the structure of high-dimensional activation features in two dimensions. UMAP is a nonlinear, stochastic dimensionality-reduction technique that aims to preserve local neighborhood relationships while maintaining coarse global structure. As a result, local geometry and exact point placement may vary across random initializations, and UMAP embeddings should be interpreted qualitatively rather than metrically. In this work, UMAP is used solely as a visualization tool to reveal large-scale organization in the learned representations; all quantitative results are computed in the original feature space. Repetition experiments assessing the robustness of the structures reported in the main paper (Section 4.2) are provided below.
D.1 Repetition Experiments on Radio Receiver
As noted above, UMAP is a stochastic dimensionality-reduction technique, and its local geometry may vary across random initializations. To assess the robustness of the observed activation structure, we repeat the UMAP embedding five times using different random seeds for DeepRx activations from layer B1-PRE. Figures 3337 show the resulting embeddings for both the baseline and proposed methodologies, colored by different metadata attributes.
Across all repetitions, the global organization of the embeddings remains stable. In particular, samples consistently form a smooth, low-dimensional manifold primarily aligned with SNR (Fig. 34), while other channel attributes such as delay spread, UE speed, and channel model induce only secondary variations (Figs. 3537). Although local neighborhoods and fine-scale layout may differ between runs–as expected for UMAP–the dominant SNR-driven structure is preserved. This confirms that the observed manifold is not an artifact of a particular UMAP initialization but reflects a robust property of the learned receiver representations.
Figure 33: UMAP embeddings of DeepRx activations from layer B1-PRE across five independent runs using the baseline (left) and proposed (right) methodology, colored by cluster partitioning. UMAP is used solely for visualization.
Figure 34: UMAP embeddings of DeepRx activations from layer B1-PRE across five independent runs using the baseline (left) and proposed (right) methodology, colored by SNR. UMAP is used solely for visualization.
Figure 35: UMAP embeddings of DeepRx activations from layer B1-PRE across five independent runs using the baseline (left) and proposed (right) methodology, colored by delay spread. UMAP is used solely for visualization.
Figure 36: UMAP embeddings of DeepRx activations from layer B1-PRE across five independent runs using the baseline (left) and proposed (right) methodology, colored by UE speed. UMAP is used solely for visualization.
Figure 37: UMAP embeddings of DeepRx activations from layer B1-PRE across five independent runs using the baseline (left) and proposed (right) methodology, colored by channel model. UMAP is used solely for visualization.
Appendix E Conditional k-NN Methodology for Out-Of-Distribution Detection
Let each training sample be represented by an activation vector $a_{i}\in\mathbb{R}^{d}$ and an associated SNR parameter $p_{i}\in\mathbb{R}$, and let the test sample and its parameter be $(a_{*},p_{*})$. In a standard $k$-nearest-neighbour (k-NN) method, only distances in activation space are considered, typically
$d_{i}=\|a_{i}-a_{*}\|,$ (34)
and the $k$ smallest distances determine the neighborhood for test sample in question. However, the relationship between training and test samples may depend on the SNR level. To capture this dependence, we introduce a conditional distance in which SNR differences influence the effective neighbour selection.
E.1 Hard Conditional Scheme
In the hard variant, SNR is used as a strict constraint. A training point is considered eligible only if its SNR is sufficiently close to the test SNR. Let $D>0$ be the maximum allowed SNR deviation. Define
$\Delta p_{i}=|p_{i}-p_{*}|.$ (35)
The effective distance is then
$d_{i}^{\text{eff}}=\begin{cases}d_{i},&\text{if }\Delta p_{i}\leq D,\\[4.0pt] \infty,&\text{otherwise}.\end{cases}$ (36)
Thus, neighbours are selected only from activation-space distances of points whose SNR is within a hard threshold of $p_{*}$, and all other points are excluded.
E.2 Soft Conditional Scheme
The hard cutoff can be replaced by a smooth penalty that grows with the SNR difference, allowing all points to contribute while favouring those with more similar SNR. Let $m(\Delta p_{i})$ be a monotone increasing penalty function with $m(0)=1$, and define the effective distance as
$d_{i}^{\text{eff}}=m(\Delta p_{i})\,d_{i}.$ (37)
In these soft schemes, the $k$ neighbours are chosen as the training points with the smallest effective distances $d_{i}^{\text{eff}}$, enabling a continuous transition between strict conditioning and unrestricted k-NN. Two practical choices for $m(\cdot)$ are power-law and exponential penalty, described below in more detail.
E.2.1 Power-law penalty.
Let $D$ denote the characteristic SNR scale and $F>1$ the penalty factor at $\Delta p_{i}=D$. With exponent $p\geq 1$,
$m(\Delta p_{i})=1+(F-1)\left(\frac{\Delta p_{i}}{D}\right)^{p}.$ (38)
This yields linear ($p=1$) or stronger ($p>1$) growth of the penalty as the SNR difference increases.
E.2.2 Exponential (Gaussian) penalty.
Define a bandwidth $h>0$ and let
$m(\Delta p_{i})=\exp\!\left(\frac{1}{2}\left(\frac{\Delta p_{i}}{h}\right)^{p}\right).$ (39)
The bandwidth $h$ may be chosen so that $m(D)$ matches a desired penalty factor $F$, leading to
$h=\frac{D}{\sqrt[p]{2\ln F}}.$ (40)
In this formulation, points with larger SNR mismatch receive exponentially inflated effective distances and thus are gradually deprioritised in the neighbour ranking.
Figure 38: OOD detection performance of k-NN-based detectors under high delay spread on layer B1-PRE as a function of the penalty exponent $p$ in (38) and (39). Mean-activation features and the cosine distance are used.
Figure 39: OOD detection performance of k-NN-based detectors under high UE speed on layer B1-PRE as a function of the penalty exponent $p$ in (38) and (39). Mean-activation features and the cosine distance are used.
E.3 Parameter Selection
The SNR tolerance $D$ is selected as the 99th percentile of absolute SNR differences observed in the ID training set, reflecting a typical operating range. The penalty factor is fixed to $F=2.0$ across experiments.
Conditional k-NN is evaluated using mean-activation features with cosine distance, as this combination yields the strongest k-NN baseline performance on average across the two OOD scenarios (Table 2). We adopt the exponential penalty, as it degrades performance less than the power-law alternative in the high delay scenario, and set the exponent to $p=1$, which provides the largest gains in the high speed scenario.
Figures 38 and 39 illustrate the sensitivity of OOD detection performance to the penalty exponent under high delay spread and high UE speed, respectively.
Appendix F Continuous Conformal P-Values Methodology for Out-Of-Distribution Detection
We briefly recall the Benjamini–Hochberg (BH)-style conformal testing rule of Algorithm 1 in magesh2023conformalpvalues. Let $\{\hat{Q}_{(i)}\}_{i=1}^{K}$ denote the ordered conformal p-values across $K$ detectors. The algorithm defines
$m=\max\left\{i:\hat{Q}_{(i)}\leq\frac{\alpha\,i}{C(K)\,K}\right\}$ (41)
and declares a sample to be Out-Of-Distribution (OOD) if
$m\geq 1.$ (42)
Equivalently, the sample is classified as OOD if there exists an index $i$ such that
$\hat{Q}_{(i)}\leq\frac{\alpha\,i}{C(K)\,K}.$ (43)
Rearranging this condition yields
$\min_{i=1,\dots,K}\left(\frac{C(K)\,K}{i}\,\hat{Q}_{(i)}\right)\leq\alpha.$ (44)
This observation allows us to convert the binary rejection rule of Algorithm 1 into a continuous score. We define the scalar
$S=\min_{i=1,\dots,K}\left(\frac{C(K)\,K}{i}\,\hat{Q}_{(i)}\right),$ (45)
which represents the smallest significance level $\alpha$ at which the sample would be declared OOD under the BH-style conformal test. While this continuous score is not introduced as a separate algorithm in prior work, it follows directly from the decision rule in magesh2023conformalpvalues.
The score $S$ provides a calibrated, continuous measure of out-of-distribution confidence and can be used to construct ROC or DET curves, enabling quantitative comparison and fusion with other OOD detectors.
Appendix G Supplemental Experimental Results on Out-Of-Distribution Detection on Radio Receiver
This section provides detailed numerical results supporting the Out-Of-Distribution (OOD) detection experiments reported in the main paper (Section 5). We organize results by analysis type–single-detector baselines, SNR-aware modeling, and detector fusion–while reporting performance separately for the three representative layers analyzed in the main text (B1-PRE, B3-POST, and OUT-POST). All results are reported for both high delay and high speed OOD scenarios, as well as their average when appropriate. Unless stated otherwise, results are averaged over five independent runs with different random seeds for DeepRx activations.
G.1 Single-Detector Baselines
We first report performance of individual OOD detectors applied to a single layer without SNR conditioning or fusion. Results for layers B1-PRE, B3-POST, and OUT-POST are reported separately in Tables 916.
G.2 Effect of SNR-Aware Modeling
We evaluate the impact of incorporating SNR information into OOD detection using two strategies: (i) conditional k-NN (hard and soft conditioning), and (ii) feature concatenation for OC-SVM and Mahalanobis detectors. Results are reported for all three layers across both OOD scenarios in Tables 1722. Soft conditional k-NN yields the largest gains under high UE speed on B1-PRE, whereas hard conditioning is more robust to increased delay spread (Figs. 3839).
G.3 Within-Layer Fusion
We evaluate the fusion of multiple detectors applied to the same layer using OC-SVM and conformal p-value aggregation. All 27 combinations of the four activation-based detectors and raw SNR are evaluated per layer:
  • B1-PRE: Tables 2426;
  • B3-POST: Tables 2729;
  • OUT-POST: Tables 3032;
and best-performing fusions compiled in Table 23.
G.4 Across-Layer Fusion
Finally, we evaluate across-layer fusion, where detector outputs from multiple layers are combined to exploit complementary information across network depth. We consider fusing raw detector outputs as well as within-layer fused scores across B1-PRE, B3-POST, OUT-POST, and SNR. All 11 combinations of the scores are evaluated per detector:
  • HDBSCAN: Tables 3436;
  • k-NN: Tables 3739;
  • OC-SVM: Tables 4042;
  • Mahalanobis: Tables 4345;
  • Within-layer: Tables 4648;
and best-performing fusions compiled in Table 33.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.9209 1.0000 0.9741 0.1812 0.9394 0.2084 0.9965 0.0200
Mean + Cosine 0.7360 1.0000 0.9931 0.0218
KDE + Euclidean 0.9011 1.0000 0.9930 0.0456 0.8628 0.4197 0.9854 0.0582
KDE + Cosine 0.8920 1.0000 0.9920 0.0526
KDE + Jensen-Shannon 0.8898 1.0000 0.9940 0.0374
Energy + Euclidean 0.8586 1.0000 0.9622 0.2538 0.8861 0.3286 0.9905 0.0458
Energy + Cosine 0.8432 1.0000 0.9675 0.2115
Table 9: OOD detection performance under high delay spread on layer B1-PRE. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.055.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.6528 1.0000 0.6670 0.9301 0.5994 0.9253 0.7351 0.8569
Mean + Cosine 0.6285 1.0000 0.7339 0.8620
KDE + Euclidean 0.6601 1.0000 0.7178 0.8199 0.6143 0.8947 0.7357 0.7864
KDE + Cosine 0.6791 1.0000 0.7164 0.8292
KDE + Jensen-Shannon 0.6715 1.0000 0.7266 0.8069
Energy + Euclidean 0.6369 1.0000 0.6691 0.9203 0.6004 0.9218 0.7179 0.8716
Energy + Cosine 0.6342 1.0000 0.6741 0.9120
Table 10: OOD detection performance under high UE speed on layer B1-PRE. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.025.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.8370 0.3543 0.9134 0.4588 0.7372 0.8597 0.9410 0.3327
Mean + Cosine 0.8807 0.2073 0.9174 0.6440
KDE + Euclidean 0.8812 1.0000 0.9323 0.3689 0.8845 0.5919 0.9389 0.3576
KDE + Cosine 0.8896 0.3904 0.9395 0.2397
KDE + Jensen-Shannon 0.8591 1.0000 0.9501 0.3023
Energy + Euclidean 0.7948 0.6573 0.9191 0.3872 0.7569 0.8255 0.9450 0.3299
Energy + Cosine 0.8201 0.3305 0.9346 0.3281
Table 11: OOD detection performance under high delay spread on layer B3-POST. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.035, except for 0.28 for HDBSCAN Energy + Euclidean.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.6955 0.5966 0.7317 0.8488 0.7387 0.8957 0.7472 0.8615
Mean + Cosine 0.7249 0.5206 0.7522 0.9138
KDE + Euclidean 0.6993 1.0000 0.7547 0.8202 0.7344 0.8538 0.7385 0.9170
KDE + Cosine 0.7210 0.6114 0.7820 0.7403
KDE + Jensen-Shannon 0.7053 1.0000 0.7704 0.7839
Energy + Euclidean 0.6864 0.7691 0.7590 0.8194 0.7279 0.9121 0.7411 0.8559
Energy + Cosine 0.6907 0.5825 0.7693 0.8039
Table 12: OOD detection performance under high UE speed on layer B3-POST. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015, except for 0.19 for HDBSCAN Energy + Euclidean.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.7662 0.4754 0.8225 0.6538 0.7379 0.8777 0.8441 0.5971
Mean + Cosine 0.8028 0.3640 0.8348 0.7789
KDE + Euclidean 0.7902 1.0000 0.8435 0.5945 0.8094 0.7229 0.8387 0.6373
KDE + Cosine 0.8053 0.5009 0.8608 0.4900
KDE + Jensen-Shannon 0.7822 1.0000 0.8603 0.5431
Energy + Euclidean 0.7406 0.7132 0.8391 0.6033 0.7424 0.8688 0.8431 0.5929
Energy + Cosine 0.7554 0.4565 0.8520 0.5660
Table 13: Average OOD detection performance over high delay spread and UE speed scenarios on layer B3-POST. Mean over 5 independent repeats with random seeds.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.5208 1.0000 0.5769 0.9495 0.5509 0.9486 0.5677 0.9359
Mean + Cosine 0.6745 1.0000 0.6929 0.8847
KDE + Euclidean 0.5692 1.0000 0.6397 0.9485 0.4978 0.9746 0.6051 0.9499
KDE + Cosine 0.5574 1.0000 0.6584 0.9484
KDE + Jensen-Shannon 0.5741 1.0000 0.6698 0.9466
Energy + Euclidean 0.5195 1.0000 0.5816 0.9495 0.5564 0.9455 0.5697 0.9481
Energy + Cosine 0.5129 1.0000 0.5900 0.9499
Table 14: OOD detection performance under high delay spread on layer OUT-POST. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.5096 1.0000 0.5303 0.9565 0.4458 0.9919 0.5204 0.9465
Mean + Cosine 0.6372 1.0000 0.6620 0.9235
KDE + Euclidean 0.5470 1.0000 0.6039 0.9536 0.4606 0.9813 0.5719 0.9505
KDE + Cosine 0.5418 1.0000 0.6182 0.9546
KDE + Jensen-Shannon 0.5499 1.0000 0.6187 0.9516
Energy + Euclidean 0.5086 1.0000 0.5279 0.9570 0.4446 0.9927 0.5147 0.9458
Energy + Cosine 0.5033 1.0000 0.5324 0.9561
Table 15: OOD detection performance under high UE speed on layer OUT-POST. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015.
Feature + Distance HDBSCAN k-NN OC-SVM Mahalanobis
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
Mean + Euclidean 0.5152 1.0000 0.5536 0.9530 0.4983 0.9703 0.5441 0.9412
Mean + Cosine 0.6558 1.0000 0.6775 0.9041
KDE + Euclidean 0.5581 1.0000 0.6218 0.9510 0.4792 0.9779 0.5885 0.9502
KDE + Cosine 0.5496 1.0000 0.6383 0.9515
KDE + Jensen-Shannon 0.5620 1.0000 0.6443 0.9491
Energy + Euclidean 0.5140 1.0000 0.5547 0.9532 0.5005 0.9691 0.5422 0.9470
Energy + Cosine 0.5081 1.0000 0.5612 0.9530
Table 16: Average OOD detection performance over high delay spread and UE speed scenarios on layer OUT-POST. Mean over 5 independent repeats with random seeds.
SNR regime k-NN OC-SVM Mahalanobis
[…]
Activations only 0.9931 0.0000 0.0218 0.0000 0.9394 0.0000 0.2084 0.0000 0.9965 0.0000 0.0200 0.0000
SNR only 0.7326 -0.2605 0.7897 0.7679 0.5000 -0.4394 1.0000 0.7916 0.5000 -0.4965 0.9589 0.9389
Both combined (hard/concat) 0.9926 -0.0005 0.0240 0.0021 0.9392 -0.0002 0.2096 0.0012 0.9696 -0.0269 0.1484 0.1285
Both combined (soft) 0.9922 -0.0009 0.0278 0.0060
Table 17: OOD detection performance under high delay spread on layer B1-PRE. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015.
SNR regime k-NN OC-SVM Mahalanobis
[…]
Activations only 0.7339 -0.0019 0.8620 0.0000 0.5994 -0.0000 0.9253 0.0001 0.7351 0.0000 0.8569 0.0000
SNR only 0.6128 -0.1230 0.9144 0.0524 0.5000 -0.0994 1.0000 0.0748 0.5000 -0.2351 0.9589 0.1020
Both combined (hard/concat) 0.7326 -0.0032 0.8660 0.0040 0.5994 0.0000 0.9252 0.0000 0.6527 -0.0823 0.8987 0.0418
Both combined (soft) 0.7358 0.0000 0.8651 0.0031
Table 18: OOD detection performance under high UE speed on layer B1-PRE. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015.
SNR regime k-NN OC-SVM Mahalanobis
[…]
Activations only 0.9174 -0.0042 0.6440 0.0433 0.7372 -0.0006 0.8597 0.0060 0.9410 0.0000 0.3327 0.0000
SNR only 0.6758 -0.2458 0.8233 0.2226 0.5000 -0.2378 1.0000 0.1463 0.5000 -0.4410 0.9589 0.6261
Both combined (hard/concat) 0.9216 0.0000 0.6007 0.0000 0.7378 0.0000 0.8537 0.0000 0.9010 -0.0400 0.3724 0.0396
Table 19: OOD detection performance under high delay spread on layer B3-POST. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015.
SNR regime k-NN OC-SVM Mahalanobis
[…]
Activations only 0.7522 -0.0089 0.9138 0.0193 0.7387 -0.0000 0.8957 0.0000 0.7472 -0.0004 0.8615 0.0463
SNR only 0.6797 -0.0814 0.8944 0.0000 0.5000 -0.2387 1.0000 0.1043 0.5000 -0.2475 0.9589 0.1437
Both combined (hard/concat) 0.7611 0.0000 0.9015 0.0071 0.7387 0.0000 0.8979 0.0022 0.7475 0.0000 0.8152 0.0000
Table 20: OOD detection performance under high UE speed on layer B3-POST. Mean over 5 independent repeats with random seeds. All standard deviations less than 0.015.
SNR regime k-NN OC-SVM Mahalanobis
[…]
Activations only 0.6929 0.0000 0.8847 0.0000 0.5509 -0.0121 0.9486 0.0362 0.5677 -0.0313 0.9359 0.0501
[…]